# Conformal Prediction for Ensembles: Improving Efficiency via Score-Based Aggregation

### v2

@@ 3.1.2 Score Quantile Threshold
To then compute $\widehat{q}$, we find $t^{*}(s)$ for each $s\in\mathcal{S}^{(2)}_{\mathcal{C}}$, defined to be $\min\{t\in\mathbb{R}:s\in\bigcap_{m=1}^{M}H(u_{m},t\widetilde{q}_{m})\}$. This can be efficiently computed as $t^{*}(s)=\max_{m=1,...,M}(u_{m}^{\top}s/\widetilde{q}_{m})$. Denoting the $\lceil(N_{\mathcal{C}_{2}}+1)(1-\alpha)\rceil$-th largest $t^{*}(s)$ as $\widehat{t}$, $\widehat{q}_{m}:=\widehat{t}\widetilde{q}_{m}$ and $\widehat{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widehat{q}_{m})$. If the tightest quantile envelope was already discovered over $\mathcal{S_{C}}^{(1)}$, this adjustment factor $\widehat{t}\approx 1$.
Importantly, while the aforementioned procedure will necessarily result in convex regions $\widehat{\mathcal{Q}}$, this does not mean the downstream prediction regions in $\mathcal{Y}$ will be convex, as discussed in Section 3.2. Such flexibility is necessary, as many studies have demonstrated the need for nonconvex prediction regions for downstream utility, such as [36, 42, 43]. However, it is unsurprising such flexibility exists, as even a single scalar score $s_{1}(x,y)$ can produce nonconvex prediction regions. We present the full algorithm in Algorithm 1.
@@
3:  $\mathcal{S_{C}}^{(1)}\cup\mathcal{S_{C}}^{(2)}\leftarrow\{(s_{k}(x_{i},y_{i}))_{k=1}^{K}\}_{i=1,N_{\mathcal{C}_{1}}+1}^{N_{\mathcal{C}_{1}},N_{\mathcal{C}_{2}}}$
4:  $\{u_{m}\leftarrow\textsc{UnifHypersphere}(K)\}_{m=1}^{M}$
5:  while $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}\notin 1-\alpha\pm\epsilon$ do
6:     $\beta\leftarrow\frac{\beta_{\mathrm{lo}}+\beta_{\mathrm{hi}}}{2}$
7:     $\left\{\widetilde{q}_{m}\leftarrow(1-\beta)\text{ emp. quantile: }\{u_{m}^{\top}s_{i}\}_{s_{i}\in S_{C}^{(1)}}\right\}_{m=1}^{M}$
8:     $\widehat{\mathcal{Q}}\leftarrow\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$
9:     if $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}>1-\alpha$ then $\beta_{\mathrm{lo}}\leftarrow\beta$
10:     else $\beta_{\mathrm{hi}}\leftarrow\beta$
11:  end while
12:  $\widehat{t}\leftarrow(1-\alpha)\text{ emp. quantile: }\{\max_{m\in[M]}\frac{u_{m}^{\top}s_{i}}{\widetilde{q}_{m}}\}_{s_{i}\in S_{C}^{(2)}}$
13:  Return $\{(u_{m},\widehat{t}\widetilde{q}_{m})\}_{m=1}^{M}$
@@ 3.2 Predict-Then-Optimize
With this generalization of the score function, a natural question is how to leverage the resulting prediction regions $\mathcal{C}(x)$. For both classification and regression, $\mathcal{C}(x)=\{y\mid s(x,y)\in\widehat{\mathcal{Q}}\}$, or $\mathcal{C}(x)=\bigcap_{m=1}^{M}\mathcal{C}_{m}(x)$ where $\mathcal{C}_{m}(x):=\{y\mid u_{m}^{\top}s(x,y)\leq\widehat{q}_{m}\}$.
In the classification setting, where $|\mathcal{Y}|\in\mathbb{N}$, explicit construction of $\mathcal{C}(x)$ is straightforward. For a fixed $x$, explicitly constructing $\mathcal{C}(x)$ can be done by iterating through $y\in\mathcal{Y}$ and checking if $s(x,y)\in\widehat{\mathcal{Q}}$ by comparing $s(x,y)$ against each one of the thresholds $\widehat{q}_{m}$ after projection.
In the case of regression, however, the prediction region cannot be explicitly constructed in the general case, since $\mathcal{Y}$ contains uncountably many elements. In fact, explicit construction is generally not of interest for downstream regression applications. Instead, prediction regions in regression settings are often simply used in their implicit form. We, therefore, focus on one particular application, namely that of [36] discussed in Section 2.3, and demonstrate the CSA prediction regions can be leveraged in their framework. Such an extension has natural applications to the settings discussed in this original study. For instance, the authors demonstrated the utility of their method in a robust traffic routing setting with $c$ being predicted traffic from a probabilistic weather model $q(C\mid X)$ for weather covariates $X$. An ensembling approach naturally emerges with having multiple predictive models, such as a $q_{2}(C\mid X)$ predicting traffic based instead on historical trends.
@@
$s_{k}(x,c)=\min_{j\in 1,...,J_{k}}\left[d_{k}\left(\widehat{c}_{kj},c\right)\right],$ (3)
where $d_{k}(\widehat{c},c)$ is a norm over the $C$ space. Notably, this framework subsumes many standard regression settings, e.g., for a deterministic predictor, one can take $q_{k}(C\mid X)=\delta(f_{k}(X))$. To compute $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, we first let $\vec{j}\in\mathcal{J}=\{j_{1},...,j_{K}\}$ be an indexing tuple, where each $j_{k}\in\{1,...,J_{k}\}$. That is, each $\vec{j}$ is a vector that “selects” one sample per predictor. Notably then, the projection $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)$ is convex in $c$, since the projection directions are all restricted to $\mathcal{S}_{+}^{K-1}$. As a result, we can compute the maximizer $c^{*}$, since
$c_{\vec{j}}^{*}$ $:=\operatorname*{arg\,max}_{c}f(w,c)$ (4)
   s.t. $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}\quad\forall m\in\{1,...,M\}$
remains a standard convex optimization problem. The final maximum can then be found by aggregation, namely $c^{*}=\operatorname*{arg\,max}_{\vec{j}\in\mathcal{J}}f(w,c_{\vec{j}}^{*})$. While $|\mathcal{J}|=\prod_{k=1}^{K}J_{k}$, we note that certain cases of ensemble prediction, such as multi-view prediction, tend to have a limited number of predictors in practice, most typically $K=2$ or $K=3$. This coupled with the fact that computing over these indices is trivially parallelizable means this approach is still computationally tractable. The full procedure is outlined in Algorithm 2.
Algorithm 2 Predict-Then-Optimize Under CSA
3:  $w^{(0)}\sim U(\mathcal{W})$
4:  for $t\in\{1,\ldots T\}$ do
5:     for $\vec{j}\in\mathcal{J}$ do $c_{\vec{j}}^{*}\leftarrow\operatorname*{arg\,max}_{c}f(w^{(t)},c)$$\mathrm{s.t.}\quad\forall m\in 1,...,M\qquad u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)\leq\widehat{q}_{m}$
8:  end for
9:  Return $w^{(T)}$
Table 1: Average coverages across tasks for $\alpha=0.10$ are shown in the top rows and average prediction set sizes in the bottom rows. Both were assessed over a batch of i.i.d. test samples (15% of the validation set from ImageNet). Standard deviations and means were computed across 10 randomized draws of the calibration and test sets.
Dataset Models ResNet VGG DenseNet VFCP $\mathcal{C}^{M}$ $\mathcal{C}^{R}$ $\mathcal{C}^{U}$ CSA
ImageNet Coverage 0.901 (0.005) 0.902 (0.003) 0.902 (0.003) 0.899 (0.004) 0.938 (0.003) 0.909 (0.004) 0.9 (0.004) 0.9 (0.003)
Size 137.004 (1.98) 136.116 (2.206) 120.096 (2.427) 46.063 (1.089) 87.337 (1.604) 82.746 (1.692) 131.856 (2.378) 34.006 (0.924)
@@ 4 Experiments
We now study CSA empirically across several tasks, demonstrating its coverage guarantees with reduced conservatism. We demonstrate improvements in an ImageNet classification task in Section 4.1, in real-data UCI regression tasks as studied in [29] in Section 4.2, and in a downstream predict-then-optimize task in Section 4.3.
To allow for fair comparison, we note that the predictors and calibration and test sets were fixed across choices of calibration procedure for each experiment. This in turn means that some care had to be taken in partitioning $\mathcal{D_{C}}=\mathcal{D_{C}}^{(1)}\cup\mathcal{D_{C}}^{(2)}$ for CSA, where an insufficiently large $\mathcal{D_{C}}^{(1)}$ would result in poor estimation of the $\alpha$-quantile envelope and hence require a large adjustment $\widehat{t}$ factor and an insufficiently large $\mathcal{D_{C}}^{(2)}$ in the classical reduced predictive efficiency from conformal prediction. For this reason, we fixed the split to be $20\%$-$80\%$ in experiments.
We herein compare against the methods presented in Section 2.4, namely the model selection of [27], the aggregation methods of [25], and the single weighted score projection (VFCP) of [40]. We specifically consider the following aggregation methods from [25]: the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches, fully described in Appendix F. Notably, these methods do not lend themselves for use in the predict-then-optimize setting, so we eliminate them from consideration therein. Similarly, VFCP can only be applied in classification settings; we, thus, do not compare to it across the regression tasks. Code will be made public upon acceptance.
@@ 4.1 Classification Tasks
We first study the predictive efficiency of the aforementioned methods on the ImageNet classification task [44]. In particular, an ensemble was constructed from three separately trained deep learning architectures, namely ResNet-50, VGG-11, and DenseNet-121. Conformalization on the individual models was performed using the standard classification score function across all approaches, namely $s(x,y)=\sum_{j=1}^{l}\widehat{f}(x)_{\pi_{j}(x)}$ where $y=\pi_{l}(x)$ and $\pi(x)$ is the permutation of $\{1,\ldots,|\mathcal{Y}|\}$ that sorts $\widehat{f}\left(x\right)$ from most to least likely. Calibration was performed using 85% of the ImageNet test set and assessment of the coverage and interval lengths on the remaining 15%, with 10 trials conducted over randomized draws of these calibration and test sets.
The results are presented in Table 1. We see that all the approaches exhibit the desired coverage guarantees. However, the CSA score approach consistently produce significantly smaller prediction regions than both the individually conformalized input views and alternate aggregation strategies.
4.2 UCI Regression Tasks
We now similarly study the predictive efficiency of CSA across a suite of regression tasks from the UCI repository [45] as studied by [29]. We specifically consider those tasks for which there are at least 1,000 samples; the complete list of tasks are provided in Appendix D. The data for each task were split with 80% for training, 10% for calibration, and 10% for testing coverage and interval lengths.
The problem setup was replicated from [29], in which four prediction methods were ensembled for each prediction task, namely an OLS model, a LASSO linear model, a random forest, and a multi-layer perceptron. A standard absolute value residual function was used as the score across all methods, namely $s(x,y)=|\widehat{f}(x)-y|$. Prediction intervals could be explicitly constructed for the $\mathcal{C}^{M}$, $\mathcal{C}^{R}$, and $\mathcal{C}^{U}$ aggregation methods. For assessing interval length for CSA, however, a discretized grid $\mathcal{G}_{Y}\subset\mathcal{Y}$ of coarseness $\Delta y$ was considered, and an interval length estimate given by $\mathcal{L}(\mathcal{C}(x))\approx\Delta y\cdot|\{y:y\in\mathcal{G}_{Y},s(x,y)\in\widehat{\mathcal{Q}}\}|$. $10$ trials were conducted for each task, randomized over subset selections of the training, calibration, and test sets. We also present an ablation, labeled “CSA (Single-Stage),” to demonstrate the two-stage calibration described in Section 3.1 is necessary to retain coverage; this single-stage approach does not split $\mathcal{S}_{C}$ and instead directly computes $\{\widehat{q}_{m}\}$ on $\mathcal{S}_{C}$ per Section 3.1.1.
A subset of the results are given in Table 2; the full set of results is deferred to Appendix D due to space limitations. As in the results presented in Section 4.1, we see that CSA retains the coverage guarantees typical of conformal prediction yet produces significantly smaller prediction intervals than both the individual models and the alternate aggregation strategies. We additionally see that the “CSA (Single-Stage)” approach fails to retain coverage, demonstrating the necessity of the two-stage calibration.

### v3

@@ 3.1.2 Score Quantile Threshold
To then compute $\widehat{q}$, we find $t^{*}(s)$ for each $s\in\mathcal{S}^{(2)}_{\mathcal{C}}$, defined to be $\min\{t\in\mathbb{R}:s\in\bigcap_{m=1}^{M}H(u_{m},t\widetilde{q}_{m})\}$. This can be efficiently computed as $t^{*}(s)=\max_{m=1,...,M}(u_{m}^{\top}s/\widetilde{q}_{m})$. Denoting the $\lceil(N_{\mathcal{C}_{2}}+1)(1-\alpha)\rceil$-th largest $t^{*}(s)$ as $\widehat{t}$, $\widehat{q}_{m}:=\widehat{t}\widetilde{q}_{m}$ and $\widehat{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widehat{q}_{m})$. If the tightest quantile envelope was already discovered over $\mathcal{S_{C}}^{(1)}$, this adjustment factor $\widehat{t}\approx 1$. Critically, such calculations can be computed efficiently in vector form. Due to space restrictions, we defer this discussion to Appendix D. We additionally there empirically validate the efficiency of the procedure under this vectorized implementation. We present the full algorithm in Algorithm 1.
Importantly, while this procedure will result in convex regions $\widehat{\mathcal{Q}}$, this does not mean the downstream prediction regions in $\mathcal{Y}$ will be convex, as discussed in Section 3.2. However, it is unsurprising such flexibility exists, as even a single scalar score $s_{1}(x,y)$ can produce nonconvex prediction regions. One additional notable property of the CSA prediction regions is that their sizes vary across $x$ even if such variability is not baked into the constituent scores. For instance, using $s_{k}(x,y):=|f_{k}(x)-y|$ with standard, scalar conformal prediction yields intervals of length $2\widehat{q}_{k}$ for any $x$, yet $|\mathcal{C}^{\mathrm{CSA}}(x)|$ even with such $\{s_{k}(x,y)\}$ will vary with $x$. This variability is desirable, as predictive uncertainty is seldom uniform across the covariate space. See Appendix E for a full illustration of this.
@@
3:  $\mathcal{S_{C}}^{(1)}\cup\mathcal{S_{C}}^{(2)}\leftarrow\{(s_{k}(x_{i},y_{i}))_{k=1}^{K}\}_{i=1,N_{\mathcal{C}_{1}}+1}^{N_{\mathcal{C}_{1}},N_{\mathcal{C}_{2}}},\quad\{u_{m}\leftarrow\textsc{UnifHypersphere}(K)\}_{m=1}^{M}$
4:  while $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}\notin 1-\alpha\pm\epsilon$ do
5:     $\beta\leftarrow(\beta_{\mathrm{lo}}+\beta_{\mathrm{hi}})/2$
6:     $\left\{\widetilde{q}_{m}\leftarrow(1-\beta)\text{ empirical quantile of }\{u_{m}^{\top}s_{i}\}_{s_{i}\in S_{C}^{(1)}}\right\}_{m=1}^{M}$
7:     $\widehat{\mathcal{Q}}\leftarrow\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$
8:     if $|\mathcal{S}^{(1)}_{\mathcal{C}}\bigcap\widehat{\mathcal{Q}}|/N_{\mathcal{C}_{1}}>1-\alpha$ then $\beta_{\mathrm{lo}}\leftarrow\beta$ else $\beta_{\mathrm{hi}}\leftarrow\beta$
9:  end while
10:  $\widehat{t}\leftarrow(1-\alpha)\text{ empirical quantile of }\{\max_{m\in[M]}(u_{m}^{\top}s_{i}/\widetilde{q}_{m})\}_{s_{i}\in S_{C}^{(2)}}$
11:  Return $\{(u_{m},\widehat{t}\widetilde{q}_{m})\}_{m=1}^{M}$
@@ 3.2 Predict-Then-Optimize
With this generalization of the score function, a natural question is how to leverage the resulting prediction regions $\mathcal{C}(x)$. For both classification and regression, $\mathcal{C}(x)=\bigcap_{m=1}^{M}\mathcal{C}_{m}(x)$ where $\mathcal{C}_{m}(x):=\{y\mid u_{m}^{\top}s(x,y)\leq\widehat{q}_{m}\}$. For classification, where $|\mathcal{Y}|\in\mathbb{N}$, explicit construction of $\mathcal{C}(x)$ is straightforward: for any $x$, explicitly constructing $\mathcal{C}(x)$ can be done by iterating through $y\in\mathcal{Y}$ and checking if $s(x,y)\in\widehat{\mathcal{Q}}$ by comparing $s(x,y)$ against each one of the thresholds $\widehat{q}_{m}$ after projection.
In the case of regression, however, the prediction region cannot be explicitly constructed in the general case, since $\mathcal{Y}$ contains uncountably many elements. In fact, explicit construction is generally not of interest for downstream regression applications. We, therefore, focus on one particular application, namely that of [36] discussed in Section 2.3, and demonstrate the CSA prediction regions can be leveraged in their framework. Such an extension has natural applications to the settings discussed in this original study. For instance, the authors demonstrated the utility of their method in a robust traffic routing setting with $c$ being predicted traffic from a probabilistic weather model $q(C\mid X)$ for weather covariates $X$. An ensembling approach naturally emerges with having multiple predictive models, such as a $q_{2}(C\mid X)$ predicting traffic based instead on historical trends.
@@
$s_{k}(x,c)=\min_{j\in 1,...,J_{k}}\left[||\widehat{c}_{kj}-c||_{2}\right].$ (2)
Notably, this framework subsumes many standard regression settings, e.g., for a deterministic predictor, one can take $q_{k}(C\mid X)=\delta(f_{k}(X))$. To compute $\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, we first let $\vec{j}\in\mathcal{J}=\{j_{1},...,j_{K}\}$ be an indexing tuple, where each $j_{k}\in\{1,...,J_{k}\}$. That is, each $\vec{j}$ is a vector that “selects” one sample per predictor. Notably then, the projection $u_{m}^{\top}s(\widehat{c}_{\vec{j}},c)$ is convex in $c$, since the projection directions are all restricted to $\mathcal{S}_{+}^{K-1}$. Thus,
[…]
remains a standard convex optimization problem. The final maximum can then be found by aggregation, namely $c^{*}=\operatorname*{arg\,max}_{\vec{j}\in\mathcal{J}}f(w,c_{\vec{j}}^{*})$. While $|\mathcal{J}|=\prod_{k=1}^{K}J_{k}$, we note that certain cases of ensemble prediction, such as multi-view prediction, tend to have a limited number of predictors in practice, most typically $K=2$ or $K=3$. This coupled with the fact that computing over these indices is trivially parallelizable means this approach is still computationally tractable. The full procedure is outlined in Algorithm 2 in Section C.1 due to space limitations.
@@ 4 Experiments
We now study CSA empirically across several tasks, demonstrating its coverage guarantees with reduced conservatism. We demonstrate improvements in an ImageNet classification task in Section 4.1, across real-data regression benchmark tasks in Section 4.2 as proposed by [42], and in a downstream predict-then-optimize task in Section 4.3. We additionally assess the robustness of CSA to imbalanced ensembles and perform an ablation study of the two-stage calibration.
We note that the predictors and calibration and test sets were fixed across choices of calibration procedure for each experiment, meaning care had to be taken in partitioning $\mathcal{D_{C}}=\mathcal{D_{C}}^{(1)}\cup\mathcal{D_{C}}^{(2)}$ for CSA, where an insufficiently large $\mathcal{D_{C}}^{(1)}$ would result in poor estimation of the $\alpha$-quantile envelope and hence require a large adjustment $\widehat{t}$ factor and an insufficiently large $\mathcal{D_{C}}^{(2)}$ in the classical reduced predictive efficiency from conformal prediction. We note the splits in each of the sections that follow.
We compare against the methods presented in Section 2.4, viz. the model selection of [27], the aggregation methods of [25], and the single weighted score projection (VFCP) of [40]. We additionally include the initial strategy discussed in Section 2.4, in which the ensemble predictor is directly conformalized, using a natural aggregate “ensemble” score, given in the following sections. From the work of [25], we consider the following methods: the standard majority-vote $\mathcal{C}^{M}$, partially randomized thresholding $\mathcal{C}^{R}$, and fully randomized thresholding $\mathcal{C}^{U}$ approaches (see Appendix L). Notably, these methods do not lend themselves for use in the predict-then-optimize setting, so we eliminate them from consideration therein. VFCP can only be applied in classification settings; we, thus, do not compare to it across the regression tasks. Code will be made public upon acceptance.